Please help me. A CUBE OF ICE 3FT BY 3FT BY 3FT. THE ICE MELTS UNTIL IT BECOMES A CUBE WHICH IS ONE HALF AS HEAVY AS THE ORIGINAL. FIND THE DIMENSION OF NEW CUBE. (DENSITY OF ICE=62LB/CUB. FT
Answer by mananth(16949)

the volume of cube is 27 cu.ft
weight = 62*27=1674 lb
half the weight =1674/2=837 lb
V=Mass/density
V=837/62
=13.5 cu.ft
volume of cube= s^3
s^3=13.5
s= cuberoot of 13.5 = 2.38 ft